# Achronal localization and representation of the causal logic from conserved current, application to massive scalar boson

### v1

@@ 7 Covariant representation of the causal logic for the massive scalar boson
The causal logic $\mathcal{C}$ is the lattice of Borel subsets of $\mathbb{R}^{4}$ which is partially ordered by set inclusion $\subset$ and which is generated and orthocomlemented by achronal separateness, i.e., the relation
@@
       The causal logic is thoroughly studied by Cegla, Jadzyk in [8] and has been studied further on. The lattice $\mathcal{C}$ enjoys some appealing physical properties similar to the ones of the lattice of abstract elementary propositions of a quantum system [8] (see also [3]): it is possible to prove that it is $\sigma$-complete irreducible, orthomodular, atomic, atomistic exactly as for a quantum lattice, but fails to satisfy the covering law and separability.
@@
       A different approach is to relate RCL to Poincaré covariant conserved (operator) density currents. See the works cited in [2]. A first concrete step in realizing a representation is done in [9] essentially showing (4.3) for a smooth conserved current with compact support.

### v2

@@ 7 Covariant representation of the causal logic for the massive scalar boson
The causal logic $\mathcal{C}$ is the lattice of Borel subsets of $\mathbb{R}^{4}$ which is partially ordered by set inclusion $\subset$ and which is generated and orthocomplemented by achronal separateness, i.e., the relation
@@
       The causal logic is thoroughly studied by Cegla, Jadczyk, Jancewicz in [9] and has been studied further on. $\mathcal{C}$ enjoys some appealing physical properties similar to the ones of the lattice of abstract elementary propositions of a quantum system [9] and see also the more recent works [3, 4]: it is possible to prove that the lattice $\mathcal{C}$ is $\sigma$-complete, irreducible, orthomodular, atomic, atomistic exactly as for a quantum lattice, but fails to satisfy the covering law and separability.
@@
       A different approach is to relate RCL to Poincaré covariant conserved (operator) density currents. See the works cited in [2] and [10]. A first concrete step in realizing a representation is done in [10] essentially showing (4.3) for a smooth conserved current with compact support.